OpenStudy (anonymous):

Question: 6 is a factor of "H". "H" is a factor 96. Thanks for viewing and fell free to answer it ! =)

OpenStudy (anonymous):

48

OpenStudy (anonymous):

24 also

OpenStudy (anonymous):

48, 24, 12

OpenStudy (anonymous):

and 12 etc

OpenStudy (anonymous):

Ok, thanks everyone!!! =)

OpenStudy (anonymous):

96/6 = 16 so H would be 6X2=12 or 6x4=24 or 6 or 6x8=48

OpenStudy (phi):

6 is a factor of "H". so we can write H= 6*m where m is the remaining factors "H" is a factor 96. so 96/H is a whole number. Using H=6m, we have 96/6m or 16/m is a whole number. 16=2*2*2*2, so m could be any combination: 16,8,4,2,1 using H= 6m, we have H= 6*(16,8,4,2,1)= 96,48,24,12,6

OpenStudy (anonymous):

ok thanks!